Reports 1-1 of 1 Clear search Modify search
AdV-SGD (FDS commissioning)
ding, vadaro - 15:56 Tuesday 05 July 2022 (56355) Print this report
Setpoints for SQZ PLLs; FIS on RR

Yesterday (Mon July 4th), I prepared the system for the SQZ measurement campaign of this week.

First there was a change in the PLLs setpoints (slow) during the week-end, for unknown reasons : 

Old one:

  • MAIN: 27029 
  • CC: 34477
  • SC: 33220

New one:

  • MAIN: 26329
  • CC: 33864
  • SC: 32035

Then the FC was aligned with both GR and IR. Once this was done I did some quick measurements of FIS with RR :

  • before 10:47 UTC : CC contrast optimization on RR
  • 10:48 UTC : scan of phase from 0 to 3.5 rad in 120s. ASQZ at 0.86 rad, sqz at 2.18 rad.
  • 10:54:00 to 11:01:00 : ASQZ phase.
  • 11:01:10 to 11:06:10 : SQZ phase. Fast unlock at 11:05:45.
  • 11:07:20 to 11:12:20 : Shot noise.

After this, the RR was removed and the system was left untouched. This morning it was found that the IR_AA was probably left on when the RR was moved in, causing IR_M3 and IR_M5 to change positions. This means that the measurements of FIS on RR had probably more losses in alignment. 

Attached is the plot of the FIS on RR. I estimate : 

SQZ level

-5.05 dB

aSQZ level

7.7 dB

Produced SQZ

8.2 dB

The losses are quite high (14%), correlating with the un-optimal alignment of the IR_M3 and IR_M5.

Images attached to this report
Comments to this report:
bawaj, ding, vardaro - 21:26 Wednesday 06 July 2022 (56368) Print this report

The change of the PLL setpoint was necessary due to the change of the reference laser frequency. We notice the jumps of the reference laser since June the 10th. The value changes by ~657 temperature steps which corresponds to ~500 MHz in the laser frequency.

PLL Main:

pll_TH0 = 26364   # till 2022 06 10
pll_TH0 = 27032   # till 2022 06 13
pll_TH0 = 26400   # till 2022 06 28
pll_TH0 = 27029   # till 2022 07 04
pll_TH0 = 26331   # current

The current value is the typical that we had since 2021.

Search Help
×

Warning

×